OpenStudy (anonymous):

this is a chemistry question.. convert 6.7 g/mL to g/in^3 using dimensional analysis

OpenStudy (anonymous):

(6.7g/ml)*(1ml/cm^3)*((2.54cm/inch)^3)= (6.7)(2.54^3)=110g/in^3